OpenStudy (mirandaregina):

Can someone please help me with my Algebra 2 homework? How do I solve the system using the substitution method? 3x + 7y = 13 Equation 1 x + 3y = -7 Equation 2

OpenStudy (shinalcantara):

3x + 7y = 13 ---- eqn1 x + 3y = -7 ------ eqn 2 from equation 2, you'll get x = -3y -7 ----eqn 3 substitute eqn 3 to eqn 1 3x + 7y = 13 3(-3y-7) + 7y = 13 -9y - 21 + 7y = 13 -2y = 13 +21 -2y = 34 y = -17 substitute the value for y to eqn 3 to get x x = -3y -7 = -3(-17) - 7 = 44